essay show all work, a community 5k run will award $50 to the winner, 35 people enter the race, and they each pay an entry fee of $20. assuming they are all equally likely to win, what is a fair price for the competition? round to the nearest cent.

thank you ahead of time for your help.

Answer by stanbon(75887)

essay show all work, a community 5k run will award $50 to the winner, 35 people enter the race, and they each pay an entry fee of $20. assuming they are all equally likely to win, what is a fair price for the competition? round to the nearest cent.
-----------------------
Probability of winning: 1/35
---
Let cost to participate be "x".
Random winnings: 50-x ; -x
----
Expected winnings = (1/35)(50-x) + (34/35)(-x)
----
A fair price means the expected winnings is zero.
-----
Solve: (1/35)(50-x) + (34/35)(-x) = 0
---
Multiply thru by 35 to get:
50-x -35x = 0
36x = 50
x = $1.39 (that is the "fair" price for entering the race)
==================
Cheers,
Stan H.